OpenStudy (amina_soneviseth):

what is the slope and y-intercept of this line? y= -2x +8 a.) slope: 8, y - intercept: -2 b.)slope: 2,y - intercept: -8 c.)slope:-1,y-intercept: 4 d.)slope:-2,y-intercept:8

OpenStudy (amina_soneviseth):

@Vocaloid please help

OpenStudy (amina_soneviseth):

@Hero help?

OpenStudy (anonymous):

\[y=mx+b\]where\[y=vertical~axis\]\[x=horizontal~axis\]\[m=slope\]\[b=y-intercept\]

OpenStudy (anonymous):

compare\[y=-2x+8\]with\[y=mx+b\]

OpenStudy (amina_soneviseth):

how? @MARC_ and thank you for the help

OpenStudy (amina_soneviseth):

ohh i get it now, thanks for the help!! @MARC_

OpenStudy (amina_soneviseth):

could you help with this one, y = x - 5

OpenStudy (amina_soneviseth):

i think its slope -5 and intercept 1 but im not sure :(

OpenStudy (anonymous):

give u a hint \[y=1x-5\]

OpenStudy (amina_soneviseth):

ohh so its slope - 1 and 5 intercept?

OpenStudy (anonymous):

how u get -1 for slope?

OpenStudy (amina_soneviseth):

umm I divided 1/5 i dont know. /(T.T)/

OpenStudy (anonymous):

use this equation y=1x-5 and try comparing with y=mx+b

OpenStudy (amina_soneviseth):

so would I multiply? 1 x 5?

OpenStudy (anonymous):

u do not need to multiply but u just need to compare only

OpenStudy (amina_soneviseth):

how would one do so?

OpenStudy (anonymous):

b gives u -5 when u compare with y=mx+b with y=1x-5 now m gives u what? (try comparing it)

OpenStudy (amina_soneviseth):

b gives me 1x?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

no y=mx+b where y=vertical axis x=horizontal axis m=slope b=y-intercept

OpenStudy (anonymous):

m=1 @amina_soneviseth

OpenStudy (anonymous):

Therefore,the slope is 1 and the y-intercept is -5

OpenStudy (anonymous):

@amina_soneviseth